An investor buys a bond for ​$5,000. The bond pays ​$207 interest every six months. After 18​ months, the investor sells the bond for ​$4,676.

Describe the types of income​ and/or loss the investor had.

The current income is ​$________.​ (Round to the nearest​ dollar.)

The investor would experience a capital gain or a capital loss or neither (choose one) in the amount of ​$________.

​( if a loss, enter a loss as a negative number rounded to the nearest​ dollar.)

The total return on this investment is ​$______​(Round to the nearest​ dollar.)
Answer #1

The current income is ​=207*3=621

The investor would experience a capital loss in the amount of 4676-5000=-$324

The total return on this investment is ​621-324=$297